OpenStudy (anonymous):

Solve for x: 5^x * 2^(x-1) = 100

OpenStudy (anonymous):

if you take the common log you can start with \[x\log(5)+(x-1)\log(2)=2\] or if you take the natural log start with \[x\ln(5)+(x-1)\ln(2)=\ln(100)\] then solve for \(x\)

OpenStudy (anonymous):

I don't know how to solve for x

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Do you factorise?